2. The valuation of preferred stock
The formula for the valuation of a share of preferred stock is P0=D/rs. In this equation, the variable D represents the (A.Coupon Payment on the share/ B.Annual dividend paid on the share/ C.Share's current Value).

Riley is considering the purchase of 350 shares of the preferred stock of Marston Manufacturing Company. The stock carries a par value of $100 per share and an annual dividend rate of 4.25%. Alternative investments of comparable risk are generating yields of 4.25%. Given this information, the per-share value of Marston’s preferred stock should be:

A. $90.00
B. $100.00
C. $125.00
D. $75.00

Riley has to postpone her purchase of Marston’s preferred shares for just over five months. By the time she is ready to invest, the return on alternative investments of comparable risk has decreased. She should expect the cost of her investment in Marston’s preferred shares to be (Less/ More) expensive.

Assume that Riley delays her investment for another few months and that when she is finally ready to make her 350-share investment in Marston, the market price of Marston’s preferred stock has changed to $135.00 per share. If she pays this price to acquire each share of Marston’s preferred stock, what rate of return will Riley earn on her investment? Remember that the shares have a par value of $100 and a dividend rate of 4.25%.

A.3.15%
B.4.10%
C.2.52%
D.2.99%
Answer #1

rate positively ..

Ans 1 Price per share = Annual dividend/required rate
Annual dividend = 100*4.25% 4.25
required rate = 4.25%
Price = 4.25/4,25% $ 100.00
ans is option B = $ 100.00
Ans 2
Market price = 135
Annual dividend = 4.25
Required rate = 4.25/135 3.15%
ans = option A) 3.15%